OpenStudy (anonymous):

Ages. Tyler is twice as old as his son. Ten years ago, Tyler was three times as old as his son. How old are they now?

OpenStudy (anonymous):

Didn't I just answer this one... deja vu

OpenStudy (anonymous):

Taylor= T, Son =S T=2*S S= 2S-10=3S OK THIS IS HARDER THAN I EXPECTED

OpenStudy (anonymous):

not quite coco T=2*S is good, but the second equation should read T-10=3(S-10)

OpenStudy (anonymous):

T=2*S T-10=3(S-10) 2*S-10=3S-30 -10+30=3S-2S 20=S T=2(S) T=40

OpenStudy (anonymous):

well done

OpenStudy (anonymous):

WOW i actuali feel like i'm learning thank you. :)